# GraspMolmo: Generalizable Task-Oriented Grasping via Large-Scale Synthetic Data Generation

### v2

@@ 3.3 GraspMolmo
Using PRISM-Train and the training data from TaskGrasp-Image (specifically, split 0 from the task split type), we create GraspMolmo by fine-tuning Molmo [15], a state-of-the-art VLM for pointing and spatial reasoning tasks, to point to grasps. We co-train on PixMo and all other Molmo training tasks in order to preserve the capability to generalize to unseen objects and settings while adapting to grasping. We use the natural language object-centric grasp description from our training data as a chain-of-thought processing step, requiring the model to output both the pixel location for the grasp and an object-centric description of that grasp. We sample 45% and 10% of our data mixture from PRISM-Train  and TaskGrasp-Image, respectively, and proportionally downweight other data sources - please refer to supplemental materials for details on data mixture and hyperparameters.
@@
[…]
@@ 4 Evaluation
We evaluate GraspMolmo and the most appropriate baseline on three distinct benchmark settings, progressively increasing in complexity and real-world applicability: a benchmark from literature with simple objects and minimal visual diversity, a synthetic held-out dataset of fully composed scenes with unseen objects, and finally, real-world transfer scenarios. The performance gap between methods widens notably as we progress from simpler to more complex evaluation scenarios, revealing fundamental differences in approach capabilities that are not apparent in basic benchmarks.
@@
Table 1: Top-1 accuracy for grasp prediction across increasingly challenging task-oriented grasping settings. Following [6], we normalize across tasks for the TaskGrasp-Image evaluation. For real-world online evaluations, we separately report the prediction success rate (was the predicted grasp correct) and the overall success rate (was the predicted grasp correct and did the robot successfully grasp the object).
4.1 Baseline
We compare our method to GraspGPT [6], the current state-of-the-art approach in Task-Oriented Grasping, with similar assumptions to ours. We also compare to Molmo [5], which is simply the base VLM we fine-tune to create GraspMolmo, in order to illustrate the utility of training on PRISM. We follow previous works [13, 6, 7] in restricting real-world evaluations to single-view RGB-D observations, and in the interest of maximizing similarity with the real-world setting, we do the same for all evaluations.
We note that GraspGPT requires a segmented point cloud of the object, and therefore depends on Segment Anything 2 (SAM2) [54] and GroundingDINO [55] to extract the object point cloud for downstream processing. Additionally, in addition to the task instruction (e.g. “give me some water”), GraspGPT also requires access to the specific object being manipulated (e.g. water bottle), and the specific action primitive being performed (e.g. handover). In the real world, such extra information may not be available, limiting its applicability. For our experiments, we use GPT-4o to infer the specific object and task primitive from the task instruction and the image, which adds both computational overhead and the potential for error propagation. In contrast to these requirements, GraspMolmo directly processes raw sensor data and freeform natural language without needing intermediate segmentation or task simplification steps.
@@ 4.2 TaskGrasp-Image
For the TaskGrasp-Image benchmark, we follow the same data presentation format as detailed in Sec. 3.2 - single-view RGB-D observations rather than fused point clouds. We evaluate on split 0 of the “task” split type from TaskGrasp, and illustrate a performance increase over all baselines in Table 1.
@@ 4.3 PRISM-Test
PRISM-Test is a fully-synthetic evaluation set constructed using the same pipeline as our training data, but with both held-out object instances (of previously seen classes) and completely novel object classes composed into complete scenes. We defer full details, including a complete object list and randomization parameters, to Section A.1. This benchmark tests generalization capabilities across novel objects and novel scenes, while maintaining the controllable diversity and large scale of synthetic data. By evaluating on completely unseen object categories, we demonstrate that GraspMolmo  learns generalizable task-grasp relationships rather than memorizing specific object-grasp pairings.
In these increasingly challenging scenarios with held-out objects and scenes, the performance differential between baselines widens. Table 1 shows our method achieves a 62.5% task-appropriate grasp success rate, while baselines drop to below 50%. This widening performance gap and lower success rates confirm our intuition that the complex scenes and tasks, along with diverse randomization, make PRISM-Test a challenging and valuable benchmark. As illustrated in Figure 4(a), we also see that PRISM-Test also correlates well with real-world performance, further proving its utility.
@@ 4.4 Real-World Transfer with PRISM-Real
Figure 3: (a) We evaluate on three real-world scenes representative of in-home use cases, with varying objects with diverse task semantics. We also demonstrate zero-shot applicability to bimanual task-oriented grasping. (b) We illustrate sample grasp outputs from GraspMolmo and baselines for the task “dump the flowers out”, where the robot must grasp the vase and turn it over, to empty out the flowers. GraspMolmo correctly grasps the vase in an optimal position to flip it.
To demonstrate real-world transfer capabilities, we test on a real robot platform. We use M2T2 [42] as the stable grasp generator, and execute the predicted grasps on a 7-DoF Franka FR2 arm. We evaluate the considered methods on 3 realistic cluttered scenes, featuring a total of 9 household objects representative of in-home applications, each with 2 associated tasks. With 3 trials per task, we perform a total of 54 trials per method. For full details about the evaluation scenes, objects, and tasks, please refer to the supplement.
Since these evaluation scenes each contain multiple objects with meaningful variation in task semantics, we can evaluate both geometric grasp success and task appropriateness in realistic settings. This focus on objects with multiple affordances and associated tasks makes for a challenging evaluation, measuring a model’s ability to reason about how the semantics of a task inform grasping behavior, meaning simply memorizing task-agnostic object semantics is insufficient. Here, GraspMolmo outperforms baselines by a considerable margin as illustrated in Table 1, attaining a 70.4% accuracy for grasp predictions, more than double that of the closest baseline. We also report a 61.1% success rate overall, including execution on the robot, whereas the closest baseline attains 31%.
We will release this benchmark as an offline evaluation set to facilitate future research in task-oriented grasping. Figure 3 presents representative images from our real-world test scenes, showcasing the complexity and diversity of the evaluation environments.
@@ 4.5 Extension to bimanual grasping
While single-arm grasping has dominated robotic manipulation research, it faces inherent limitations for tasks requiring coordinated forces. Task-semantic bimanual grasping addresses these constraints by enabling opposing force application—essential for activities like unscrewing a water bottle cap, folding clothes, or lifting large objects. Our preliminary experiments in this area show that GraspMolmo preserves single-arm generalization capabilities while enabling task-semantic reasoning through prompt modifications. A qualitative sample may be seem in Fig.3A - while limited, this demonstrates that GraspMolmo supports new directions in robotic manipulation.
@@ 5 Conclusion
Our work demonstrates significant advances in robotic grasping capabilities, with performance improvement on nontrival task semantics and zero-shot generalization to cluttered real scenes. By developing a system that understands not just what an object is, but how it should be manipulated for specific intended uses, we aim to advance robotic manipulation beyond simple pick-and-place operations. Our approach demonstrates superior performance on realistic scenarios, successfully navigating cluttered environments and complex language instructions like “mince some garlic” rather than the constrained “grasp [noun] to [verb]” paradigm of previous work. The zero-shot generalization to real scenes validates our synthetic training methodology, while our zero-shot success with bimanual grasping showcases the flexibility of our task-semantic understanding. We will release GraspMolmo, PRISM-Train, PRISM-Test, and PRISM-Real, as well as our code for generating synthetic data to accelerate expansion in this direction. These contributions establish a new foundation for robotic manipulation by demonstrating a generalizable approach that functions effectively in cluttered environments while showing promising potential for zero-shot extension to bimanual tasks—a crucial step toward truly effective manipulation in unstructured real-world settings.
@@ A.1 PRISM Scene Generation
Light azimuth (rad) $[0,2\pi]$
Light inclination (rad) $[0,\pi/3]$
@@ A.7 Training GraspMolmo
A.8 Real-Robot Evaluation
In Section 4 we test multiple TOG models on multiple challenging semantic grasping tasks in realistic cluttered scenes. We enumerate the manipulated objects and the associated tasks in Table 5. Notably, we select the tasks such that the robot must grasp the object in different ways. This ensures that we test a model’s ability to understand task-level semantics (where to grasp an object for a particular task) rather than simply memorizing object-level semantics (where to grasp an object independent of task).

### v3

@@ 3.3 GraspMolmo
Using PRISM-Train and the training data from TaskGrasp-Image (specifically, split 0 from the task split type), we create GraspMolmo by fine-tuning Molmo [5], a state-of-the-art VLM for pointing and spatial reasoning tasks, to point to grasps. We co-train on PixMo and all other Molmo training tasks in order to preserve the capability to generalize to unseen objects and settings while adapting to grasping. We use the natural language object-centric grasp description from our training data as a chain-of-thought processing step, requiring the model to output both the pixel location for the grasp and an object-centric description of that grasp. We sample 45% and 10% of our data mixture from PRISM-Train  and TaskGrasp-Image, respectively, and proportionally downweight other data sources. We defer details on data mixture and hyperparameters to Appendix A.7, and prompting details to Appendix A.8.
@@
GraspMolmo outputs points on the image plane, which must be matched to a candidate grasp predicted by a stable grasp generator. To do so, we map each grasp to a point on the image, and select the grasp with the closest corresponding point to Molmo’s prediction. Concretely, given a set of candidate grasps $\mathcal{G}\subset\text{SE(3)}$, a function $f\colon\mathcal{G}\to\mathbb{R}^{2}$ that maps a grasp to a pixel coordinate, and the output $p\in\mathbb{R}^{2}$ from Molmo, we select the grasp $\hat{g}\coloneqq\arg\min_{g\in\mathcal{G}}\|f(g)-p\|$. Additional details on this process are given in Appendix A.8.
@@ 4 Evaluation
We evaluate GraspMolmo and baselines on three distinct benchmark settings: a benchmark from literature with simple objects and minimal visual diversity, a synthetic held-out dataset of fully composed scenes with unseen objects, and real-world transfer scenarios. The performance gap between methods widens notably as we progress from simpler to more complex evaluations, revealing fundamental differences in approach capabilities that are not apparent in basic benchmarks.
@@
Table 1: Top-1 accuracy for grasp prediction across increasingly challenging task-oriented grasping settings. For real-world online evaluations, we separately report the prediction success rate (was the predicted grasp correct) and the overall success rate (was the predicted grasp correct and did the robot successfully grasp the object).
4.1 Baselines
We compare our method to GraspGPT [6], the current state-of-the-art approach in Task-Oriented Grasping, with similar assumptions to ours. We use the pretrained mode_t_split_0 checkpoint of GraspGPT without modification. We also compare to Molmo [5], which is simply the base VLM we fine-tune to create GraspMolmo, in order to illustrate the utility of training on PRISM. Finally, we also include a naive random baseline that uniformly samples a grasp from the candidates, underlining the importance of task-oriented reasoning. We follow previous works [2, 6, 7] in restricting real-world evaluations to single-view RGB-D observations, and in the interest of maximizing similarity with the real-world setting, we do the same for all evaluations.
We note that GraspGPT requires a segmented point cloud of the object, and therefore depends on Segment Anything 2 (SAM2) [52] and GroundingDINO [53] to extract the object point cloud for downstream processing. Additionally, in addition to the task instruction (e.g. “give me some water”), GraspGPT also requires access to the specific object being manipulated (e.g. water bottle), and the specific action primitive being performed (e.g. handover). In the real world, such extra information may not be available, limiting its applicability. For GraspGPT in our setting, we use GPT-4o to infer the specific object and task primitive from the task instruction and the image, adding both computational overhead and the potential for error propagation. In contrast to these requirements, GraspMolmo directly processes raw sensor data and freeform natural language without needing intermediate segmentation or task simplification steps.
@@ 4.2 TaskGrasp-Image
For the TaskGrasp-Image benchmark, we follow the same data presentation format as detailed in Sec. 3.2 - single-view RGB-D observations rather than fused point clouds. We evaluate on split 0 of the “task” split type from TaskGrasp, and illustrate a performance increase over all baselines in Table 1. Following [6], we normalize across tasks for this evaluation.
@@ 4.3 PRISM-Test
PRISM-Test is a synthetic evaluation set constructed using the same pipeline as our training data, but with both held-out object instances (of seen classes) and completely novel object classes. We defer full details, including a complete object list and randomization parameters, to Appendix A.1. This benchmark tests generalization capabilities across novel objects and novel scenes, while maintaining the controllable diversity and large scale of synthetic data. By evaluating on completely unseen object categories, we demonstrate that GraspMolmo learns generalizable task-grasp relationships rather than memorizing specific object-grasp pairings.
In these challenging scenarios with held-out objects and scenes, Table 1 shows our method achieves a 62.5% success rate, while baselines drop below 50%. This widening performance gap and lower success rates confirm our intuition that the complex scenes and tasks, along with diverse randomization, make PRISM-Test a challenging and valuable benchmark. As illustrated in Figure 4(a), we also see that PRISM-Test also correlates well with real-world performance, further proving its utility.
@@ 4.4 Real-World Transfer with PRISM-Real
Figure 3: (a) We evaluate on three real-world scenes representative of in-home use cases, with varying objects with diverse task semantics. We also demonstrate zero-shot applicability to bimanual task-oriented grasping. Some task instructions have been shortened for brevity. (b) We illustrate sample grasp outputs from GraspMolmo and baselines for the task “dump the flowers out”, where the robot must grasp the vase and turn it over, to empty out the flowers. GraspMolmo correctly grasps the vase in an optimal position to flip it.
To demonstrate real-world transfer capabilities, we test on a real robot platform. We use M2T2 [40] as the stable grasp generator, and execute the predicted grasps on a 7-DoF Franka FR3 arm. We evaluate the considered methods on 3 realistic scenes, featuring a total of 9 household objects representative of in-home applications, each with 2 associated tasks. With 3 trials per task, we perform a total of 54 trials per method. Evaluation scenes and sample tasks are pictured in Fig. 3, showcasing the complexity and diversity of our evaluation. For full details about the evaluation scenes, objects, and tasks, please refer to Appendix A.9.
Not only do our evaluation scenes contain multiple objects with meaningful variation in task semantics, but each object has multiple affordances and associated tasks. This measures a model’s ability to reason about how task semantics inform grasping behavior, going beyond simple memorization of task-agnostic affordances. We see in Table 1 that GraspMolmo outperforms baselines by a considerable margin, attaining a 70.4% accuracy for grasp predictions, double that of the closest baseline. We also report a 61.1% success rate overall, including execution on the robot, whereas the closest baseline attains 33.3%. A paired t-test confirms that our improvements over Molmo and GraspGPT are statistically significant at $p<0.05$, validating our real-world evaluation.
GraspMolmo’s sim-to-real transfer is largely attributable to a few key design choices. Firstly, GraspMolmo is co-trained with real images, both from Pixmo [5] and real grasping data from TaskGrasp-Image. As shown in Table 4, real images comprise 55% of training data, helping to combat overfitting to unrealistic simulation renderings. Additionally, PRISM-Test tends to be more adversarial than real-world deployment scenarios, due to aggressive randomization of camera pose, object configuration, and more, which aids transfer.
@@ 4.5 Extension to bimanual grasping
While single-arm grasping has dominated robotic manipulation research, it faces inherent limitations for tasks requiring coordinated forces. Task-semantic bimanual grasping addresses these constraints by enabling opposing force application—essential for activities like unscrewing a water bottle cap, folding clothes, or lifting large objects. Our preliminary experiments in this area show that GraspMolmo preserves single-arm generalization capabilities while enabling task-semantic reasoning through prompt modifications. A qualitative sample may be seem in Fig. 3A - while limited, this demonstrates that GraspMolmo supports new directions in robotic manipulation. We defer additional details to Appendix A.10.
@@ 5 Conclusion
Our work demonstrates significant advances in robotic grasping capabilities, with performance improvement on nontrival task semantics and zero-shot generalization to complex real scenes. By developing a system that understands not just what an object is, but how it should be manipulated for specific intended uses, we aim to advance robotic manipulation beyond simple pick-and-place operations. Our approach demonstrates superior performance on realistic scenarios, successfully navigating busy environments and complex language instructions like “mince some garlic” rather than the constrained “grasp [noun] to [verb]” paradigm of previous work. The zero-shot generalization to real scenes validates our synthetic training methodology, while our zero-shot success with bimanual grasping showcases the flexibility of our task-semantic understanding. We release GraspMolmo, PRISM-Train, PRISM-Test, and PRISM-Real, as well as our code for generating synthetic data to accelerate expansion in this direction. These contributions establish a new foundation for robotic manipulation by demonstrating a generalizable approach that functions effectively in complex and realistic environments while showing promising potential for zero-shot extension to bimanual tasks—a crucial step toward truly effective manipulation in unstructured real-world settings.
@@ A.1 PRISM Scene Generation
Light azimuth (rad) $0,2\pi]$
Light inclination (rad) $0,\pi/3]$
@@ A.7 Training GraspMolmo
Since GraspMolmo outputs a point on the image plane, we must supervise the training with 2D points rather than 6-DOF grasps. To map each grasp in PRISM to a point on the image, we perform the following process. For each sampled grasp on each object, as described in Section A.2, we raycast from the grasp TCP along the grasp direction onto the object mesh. The resulting point on the object mesh is the “grasp point” corresponding to the grasp, which is used to supervise GraspMolmo’s point prediction. In a very small number of cases, this raycast may not intersect the object mesh (e.g. if the grasp is off-center of the mesh), in which case we simply select the point on the mesh closest to the grasp TCP.
@@
A.8 GraspMolmo and Molmo Inference Details
When performing inference with Molmo and GraspMolmo, specific phrasing of the prompt may affect performance. In both training and evaluation, we use the following prompt for GraspMolmo: “Point to the grasp that would accomplish the following task: <task>”. For Molmo evaluation, we use the following prompt: “Point to where I should grasp to accomplish the following task: <task>”, which we found to produce better results for Molmo than the former. In both prompt templates, <task> should be replaced with the relevant task command, e.g. “Hand me the knife safely”.
As described in Section 3.3, GraspMolmo (and therefore Molmo as well) outputs a point $p\in\mathbb{R}^{2}$ on the image, which must be mapped to a candidate grasp to be executed. Naively, one could consider back-projecting $p$ to 3D space using depth information and directly choosing the closest grasp in cartesian space. However, this is suboptimal due to (a) depth information being incomplete and therefore not necessarily available for every pixel, and (b) small errors in pixel space translating to large errors in 3D space at object boundaries, resulting in nearby points in pixel space being very far apart in cartesian space.
To address these problems, we instead transform every candidate grasp to pixel space and perform the closest-point matching there. While this does not completely solve the problem of disparate grasps being mapped to similar points, it at least does not incorrectly discard grasps on the object being pointed to, thereby increasing recall. Directly projecting every candidate grasp to pixel space is suboptimal, since the VLM outputs the point on the surface of the object to be grasped, which may not coincide with the origin of the grasp (i.e. the TCP), or any specific point in the grasp frame, so we opt for a slightly more sophisticated approach. Concretely, given the scene point cloud, for each grasp candidate we find every point within the (non-axis aligned) box bounded by the gripper fingers, and of those, select the one closest to the gripper origin. We then project this point to the image plane to get the pixel point corresponding to the grasp candidate. Finally, finding the closest such point to the VLM’s output yields the predicted grasp.
A.9 Real-Robot Evaluation
In Section 4 we test multiple TOG models on multiple challenging semantic grasping tasks in realistic and busy scenes. We enumerate the manipulated objects and the associated tasks in Table 5. Notably, we select the tasks such that the robot must grasp the object in different ways. This ensures that we test a model’s ability to understand task-level semantics (where to grasp an object for a particular task) rather than simply memorizing object-level semantics (where to grasp an object independent of task).
As noted in Section 6, GraspMolmo has some limitations that could be addressed in future work. In our evaluations, these limitations manifest as some common failure modes. These modes include (a) pointing to the wrong part of the correct object, (b) pointing to the wrong object, and (c) pointing to the correct location, but matching the point to the incorrect grasp. Respectively, these modes account for 62%, 24%, and 5% of failures on PRISM-Real.
@@ A.9 Real-Robot Evaluation
A.10 Bimanual Task-Oriented Grasping Details
In Section 4.5 we provide a preliminary evaluation of GraspMolmo on a bimanual platform, allowing it to accomplish a wider variety of tasks with more complexity compared to single-arm TOG systems. This is provided as a showcase of possible capabilities, but more thorough methodology and evaluation are beyond the scope of this work. To get bimanual grasps to open the water bottle, as pictured in Figure 3, we query GraspMolmo twice with the tasks “hold the water bottle” and “open the water bottle”. Future work can extend and improve upon task decomposition to ameliorate bimanual TOG.